# Smoothing Variances Across Time: Adaptive Stochastic Volatility

### v4

@@ Abstract
We introduce a novel Bayesian framework for estimating time-varying volatility by extending the Random Walk Stochastic Volatility (RWSV) model with Dynamic Shrinkage Processes (DSP) in log-variances. Unlike the classical Stochastic Volatility (SV) or GARCH-type models with restrictive parametric stationarity assumptions, our proposed Adaptive Stochastic Volatility (ASV) model provides smooth yet dynamically adaptive estimates of evolving volatility and its uncertainty. We further enhance the model by incorporating a nugget effect, allowing it to flexibly capture small-scale variability while preserving smoothness elsewhere. We derive the theoretical properties of the global-local shrinkage prior DSP. Through simulation studies, we show that ASV exhibits remarkable misspecification resilience and low prediction error across various data-generating processes. Furthermore, ASV’s capacity to yield locally smooth and interpretable estimates facilitates a clearer understanding of the underlying patterns and trends in volatility. As an extension, we develop the Bayesian Trend Filter with ASV (BTF-ASV) which allows joint modeling of the mean and volatility with abrupt changes. Finally, our proposed models are applied to time series data from finance, econometrics, and environmental science, highlighting their flexibility and broad applicability.
@@ 1 Introduction
Volatility of a time series quantifies deviations from the mean and is fundamental to understanding the underlying data-generating process. In finance, volatility estimation is essential for pricing, risk assessment, and asset management (Bollerslev, 1986; Hull and White, 1987). In epidemiology, it aids in early outbreak detection and forecasting new cases of diseases (Sarkar and Chatterjee, 2017; Achcar et al., 2020; Kostoulas et al., 2021). Climate science measures volatility to study phenomena like tornadoes, droughts, and rainfall (Tippett, 2014; Modarres and Ouarda, 2014; Mehdizadeh et al., 2017). In engineering, volatility helps predict mechanical failures (Pham and Yang, 2010; Ma et al., 2017), while in hydrology, it informs understanding of streamflow variations (Otache, 2012; Wang et al., 2005, 2023a, 2023b).
@@
Figure 1: Figures 1(a), 1(b) and 1(c) display the S&P 500 index from 2012-01-01 to 2021-12-31, daily total electricity demand (in gigawatt hours) in New York State from 2015-07-01 to 2024-06-30, and daily bike rental counts via Capital Bikeshare by non-members in Washington metropolitan area from 2012-01-01 to 2017-12-31, respectively. Figures 1(d), 1(e) and 1(f) show the corresponding mean-centered series: weekly log returns for the S&P 500 index with the average log return subtracted, and series for electricity demand and bike rentals mean-centered using smoothing spline, respectively. The S&P 500 index data are obtained using the quantmod package in R (Ryan and Ulrich, 2024) from Yahoo Finance (Finance, 2025), electricity demand data from the New York Independent System Operator (NYISO) (EIA, 2024), and bike rental data from Capital Bikeshare (Bikeshare, 2024).
@@
Classical volatility models include Autoregressive Conditional Heteroskedasticity (ARCH) (Engle, 1982), Generalized ARCH (GARCH) (Bollerslev, 1986), and Stochastic Volatility (SV) (Hull and White, 1987; Taylor, 2008; Melino and Turnbull, 1990). They all make the stationarity assumption that the unconditional variance is constant. However, a growing body of econometric literature points to evolving volatility patterns in stock markets around the world (Chou, 1988; French et al., 1987; Poon and Taylor, 1992; So et al., 1997; Su and Wang, 2020; Nishino and Kakamu, 2015). Moreover, the stationarity assumption is often unrealistic in fields like climate science and epidemiology, due to climate change and seasonal or outbreak-driven dynamics (Tippett, 2014; Kostoulas et al., 2021).
@@
Figure 1 illustrates evolving volatility patterns in three real-world datasets: weekly log-returns of the S&P 500 index, daily electricity demand in New York State, and daily bike rental counts via Capital Bikeshare. All three exhibit nonstationary volatility patterns. In Figures 1(a) and 1(d), persistent periods of high and low volatility reflect changing market conditions. In Figures 1(b), 1(e), 1(c) and 1(f), both electricity and bike data show seasonal cycles in both mean and volatility, with elevated volatility during high-demand periods.
@@
Existing approaches address such patterns by incorporating hidden Markov switching (HMS) structure into existing SV and GARCH models (Brunetti et al., 2008; Ardia et al., 2018; Hamilton and Susmel, 1994; Cai, 1994; Bauwens et al., 2010; Gray, 1996a; So et al., 1998; Hwang et al., 2004). However, HMS models suffer from three major limitations: they (1) require specification of the number of regimes, which is typically unknown; (2) are computationally intensive as pointed out in Gray (1996b) and Billio et al. (2016); and (3) impose a discrete regime-switching structure, which may not capture ever evolving volatility dynamics. For example, Figure 1(f) reveals a seasonal increase in volatility for bike rentals, with the magnitude of these fluctuations growing over time. For example, volatility in 2012 is substantially lower than in 2017. Such a pattern is difficult to reconcile with a model based on a fixed number of regimes, as it reflects a continuous evolution of volatility rather than discrete shifts.
@@
We introduce the Adaptive Stochastic Volatility (ASV) model as an extension of the Random Walk Stochastic Volatility (RWSV) model (Ruiz, 1994; Harvey et al., 1994). RWSV assumes the log-variance of the observed process follows a Gaussian random walk with a constant variance. In contrast, ASV introduces two significant modifications to RWSV: it incorporates a time-varying variance for the log-variance increments and adopts a global-local shrinkage prior for adaptability. For the shrinkage prior, we specifically highlight a class of prior distributions called Dynamic Shrinkage Processes (DSP) by Kowal et al. (2019), as it provides a flexible framework for adaptive shrinkage for time series. To further improve robustness, we also consider a nugget variant of ASV, which adds an additional Gaussian noise to capture local fluctuations not explained by the smooth latent process.
@@
The standout feature of our proposed model is its local adaptability, effectively estimating volatility in the presence of both gradual and abrupt changes in the volatility process with notable smoothness in between. It allows robustness against model misspecification, resulting in low prediction error across diverse data-generating processes (DGPs). In addition, ASV’s smooth volatility estimates provide clearer insights into the underlying patterns and trends. The model also naturally highlights periods of heightened volatility variation, offering additional interpretability in identifying structural shifts.
@@
The remainder of this paper is organized as follows. Section 2 first introduces DSP and its theoretical properties, followed by a description of our proposed model (ASV) and its associated parameters. Section 3 provides an overview of the data augmentation techniques for efficient Gibbs sampling scheme; full conditional distributions are detailed in the supplementary material. Simulation study results are summarized in Section 4. Section 5 presents empirical analyses on three real-world datasets: the S&P 500 index, daily electricity usage in New York, and bike rental counts. In Section 6, the Bayesian Trend Filter with ASV (BTF-ASV), which jointly estimates the time-varying means and the variances, is proposed as an extension and applied to the sunspot number data.
@@ 2.1 The Random Walk Stochastic Volatility Model
ASV may be viewed as an extension of RWSV (Ruiz, 1994; Harvey et al., 1994). Consider a zero mean process with $T$ observations, $\{y_{t}\}_{t=1}^{T}$, and its log-variance term $\{h_{t}\}_{t=1}^{T}$. SV assumes the observed process $y_{t}$ to follow a normal distribution with its log-variance term, $h_{t}$, following the lag order 1 autoregression. RWSV is a special case of SV in which the autoregressive coefficient is fixed at 1. Thus, the model is defined as:
@@
In RWSV, $h_{t}$ is governed by a time-invariant variance term $\sigma_{h}^{2}$, which determines the degree of variation between successive log-variances. A large $\sigma_{h}^{2}$ implies high probability of significant changes in the $h_{t}$ process, whereas a small $\sigma_{h}^{2}$ indicates high probability of minimal to no changes in $h_{t}$. In a frequentist framework, $\sigma_{h}$ is treated as a fixed parameter, often estimated via quasi-maximum likelihood (Ruiz, 1994; Harvey et al., 1994). A more recent study by Nishino and Kakamu (2015) employs a Bayesian framework for estimation with a non-informative inverse-gamma priors for $\pi(\sigma^{2}_{h})$.
@@ 2.2 Stochastic Volatility with Local Only Shrinkage
RWSV may fit poorly when the time series exhibits both significant and minimal changes, as it assumes a time-invariant innovation variance $\sigma^{2}_{h}$. Ideally, we would prefer $\sigma^{2}_{h}$ to adapt over time, remaining small during stable periods and increasing during periods of abrupt changes in $h_{t}$. A natural approach to address this limitation is to introduce sparsity in the variance process. Under the frequentist framework, sparsity in $\Delta h_{t}$ may be attained by imposing a penalty, such as the $\ell_{1}$ penalty used in LASSO (Tibshirani, 1996). In a Bayesian setting, near sparsity can be imposed via heavy-tailed priors on the variance parameter. Motivated by the Bayesian LASSO (Park and Casella, 2008) and the scale-mixture Gaussian representation of the Laplace distribution (West, 1987), we propose an intermediate model, RWSV with Bayesian LASSO (RWSV-BL):
$y_{t}=\exp\{h_{t}\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (2)
$[\Delta h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma^{2}_{h,t}|\Lambda)$ $[\sigma^{2}_{h,t}]\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Exp}\{(2\Lambda^{2})^{-1}\},$
where $\Lambda>0$ is a hyperparameter controlling the prior concentration. Following Park and Casella (2008), we place a Gamma prior on $\Lambda^{2}$ for full Bayesian inference. Under RWSV, $\sigma^{2}_{h,t}$ induces shrinkage by encouraging most increments to be small, while still allowing occasional large deviations.
@@ 2.3 Dynamic Shrinkage Processes and their Properties
Before introducing the proposed model, we first describe a class of global-local shrinkage priors known as DSP, a central component in our formulation. Global-local shrinkage priors have been widely studied for Gaussian observations, particularly in high-dimensional regression. Notable examples include the horseshoe prior (Carvalho et al., 2010), the horseshoe+ prior (Bhadra et al., 2015), and the triple gamma prior (Cadonna et al., 2019). Among these, DSP offers a particularly flexible formulation that generalizes several existing priors and naturally incorporates temporal dependence.
@@
DSP has shown versatility across various applications. For example, Wu et al. (2024a) integrate DSP into a Bayesian dynamic linear model to estimate change points and score outliers. Schafer and Matteson (2024) apply it to construct the negative binomial Bayesian trend filter (NB-BTF) for adaptively smoothing integer-valued time series. Additionally, Wu et al. (2024b) combine the Bayesian trend filter with DSP and a machine learning-based regularization method to effectively distinguish micro-level drifts from macro-level shifts.
@@
DSP serves as a prior for time-varying log-variance parameters. Define $v_{t}:=\log(\sigma^{2}_{h,t})$, where $v_{t}|\mu,\phi\sim DSP(a,b,\mu,\phi)$. DSP imposes an autoregressive structure on $\{v_{t}\}_{t=1}^{T}$, allowing for locally adaptive shrinkage through heavy-tailed innovations around a persistent mean process:
$v_{t}:=\log(\sigma^{2}_{h,t})=\mu+\phi(v_{t-1}-\mu)+\eta_{t}$ $[\eta_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}Z(a,b,0,1),$ (3)
where $Z$-distribution has the following density function (Barndorff-Nielsen et al., 1982):
$f(z|a,b,\mu_{z},\sigma_{z})=(\sigma_{z}\mathrm{B}(a,b))^{-1}\bigg{(}\exp\bigg{\{}\frac{z-\mu_{z}}{\sigma_{z}}\bigg{\}}\bigg{)}^{a}\bigg{(}1+\exp\bigg{\{}\frac{z-\mu_{z}}{\sigma_{z}}\bigg{)}\bigg{\}}^{-(a+b)}.$
@@
For fully Bayesian inference, we place weakly informative priors on $\mu$ and $\phi$, and describe them in detail in Section 3. Under this framework, the global parameter is defined as $\tau=\exp\{\mu(1-\phi)/2\}$ and the local shrinkage parameter is defined as $\lambda_{t}=\exp\{(\phi v_{t-1}+\eta_{t})/2\}$. The autoregressive coefficient $|\phi|<1$ controls temporal dependence in local shrinkage. When $\phi$ is close to 1, the process exhibits strong positive dependence: strong shrinkage tends to be followed by continued strong shrinkage, and weak shrinkage by weak shrinkage. When $\phi$ is close to -1, the process favors alternation between strong and weak shrinkage across time.
@@
The $Z$-distribution on the local scale parameter $\eta_{t}$ governs the degree of local shrinkage at each time $t$. In our simulation study (Section 4) and empirical study (Section 5), we consider two specific instances of DSP. When $a=b=1/2$ and the autoregressive coefficient $\phi=0$, the prior reduces to the well-known horseshoe prior (HS) (Carvalho et al., 2010), since $\lambda_{t}=\exp\{\eta_{t}/2\}\sim C^{+}(0,1)$. When $\phi$ is estimated from the data with $a=b=1/2$, the prior induces temporal dependence in shrinkage and is referred to as the dynamic horseshoe (DHS) prior (Kowal et al., 2019). Other special cases of DSP include the Strawderman–Berger prior when $a=1/2$ and $b=1$ (Strawderman, 1971), the normal–exponential–gamma prior when $a=1$ and $b>2$ (Griffin and Brown, 2005), and the Jeffreys prior with $a=b\rightarrow 0$ (Figueiredo, 2003).
@@
While DSP has an appealing formulation and a clear connection to well-known priors, its theoretical properties remain incomplete. In particular, DSP distinguishes itself from other global-local shrinkage priors by introducing temporal dependence in the local shrinkage parameter $\lambda_{t}$. Due to the time dependence, the original work by Kowal et al. (2019) focuses on the conditional behavior while the properties of the marginal distribution remain unexamined. We derive new results on the stationary distribution of DSP.
@@
Define $z_{t,h}:=\phi^{h}\eta_{t-h}$. Infinite-order moving average representation of $v_{t}$ results in $v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t}=\mu+\sum_{h=0}^{\infty}z_{t,h}$, with $\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$. As shown in Barndorff-Nielsen et al. (1982), $z_{t,h}$ is a scaled hyperbolic secant distribution, with $\mathbb{E}(z_{t,h}|\phi)=0$ and $\mathrm{Var}(z_{t,h}|\phi)=(|\phi|^{h}\pi)^{2}$. Consider $z_{t}:=\sum_{h=0}^{\infty}z_{t,h}.$
@@
$z_{t}$ converges almost surely if and only if $|\phi|<1.$
@@
This directly follows from Kolmogorov’s three series theorem. Note that the variance $\sum_{h=0}^{\infty}\mathrm{Var}(z_{t,h}|\phi)$ converges if and only if $|\phi|<1$. Exact derivation is explored in Appendix A.1. ∎
@@
When $|\phi|<1$,
$\mathbb{E}(z_{t}|\phi)=\mathbb{E}(\sum_{h=0}^{\infty}z_{t,h}|\phi)=\sum_{h=0}^{\infty}\mathbb{E}(z_{t,h}|\phi)=0.$
$\mathrm{Var}(z_{t}|\phi)=\mathrm{Var}(\sum_{h=0}^{\infty}z_{h}|\phi)=\sum_{h=0}^{\infty}\mathrm{Var}(z_{h}|\phi)=\sum_{h=0}^{\infty}\pi^{2}\phi^{2h}=\frac{\pi^{2}}{1-\phi^{2}}.$
Characteristic function of $z_{h,t}$ is $\mathrm{sech}(\pi\phi^{h}t)$. Thus, the characteristic function, $g(t)$, for $\sum_{h=0}^{\infty}z_{h,t},$ would be an infinite product of the characteristic function of $z_{h,t}$.
$g(t)=\prod_{h=0}^{\infty}sech(\pi\phi^{h}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
Similarly, for the moment generating function $M(t)$
$M(t)=\prod_{h=0}^{\infty}\mathrm{sec}(\pi\phi^{h}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
@@
If $|\phi|=0.5$, $\eta_{t-h}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$, and $z_{t,h}:=\phi^{h}\eta_{t-h}$, then $z_{t}:=\sum_{h=0}^{\infty}z_{t,h}\stackrel{{\scriptstyle a.s}}{{\to}}Logistic(0,2)$.

### v5

@@ Abstract
We introduce a novel Bayesian framework for estimating time-varying volatility by extending the Random Walk Stochastic Volatility (RWSV) model with Dynamic Shrinkage Processes (DSP) in log-variances. Unlike the classical Stochastic Volatility (SV) or GARCH-type models with restrictive parametric stationarity assumptions, our proposed Adaptive Stochastic Volatility (ASV) model provides smooth yet dynamically adaptive estimates of evolving volatility and its uncertainty. We further enhance the model by incorporating a nugget effect, allowing it to flexibly capture small-scale variability while preserving smoothness elsewhere. We derive the theoretical properties of the global-local shrinkage prior DSP. Simulation studies demonstrate that ASV is highly robust to misspecification, consistently recovering the latent volatility structure across a wide range of data-generating processes. Furthermore, ASV’s capacity to yield locally smooth and interpretable estimates facilitates a clearer understanding of the underlying patterns and trends in volatility. As an extension, we develop the Bayesian Trend Filter with ASV (BTF-ASV) which allows joint modeling of the mean and volatility with abrupt changes. Finally, our proposed models are applied to time series data from finance, econometrics, and environmental science, highlighting their flexibility and broad applicability.
@@ 1 Introduction
Volatility of a time series quantifies deviations from the mean and is fundamental to understanding the underlying data-generating process. In finance, volatility estimation is essential for pricing, risk assessment, and asset management (GARCH2; SVop). In epidemiology, it aids in early outbreak detection and forecasting of new cases of diseases (Epidem2; Epidem; epidem3). Climate science measures volatility to study phenomena such as tornadoes, droughts, and rainfall (climate1; climate2; climate3). In engineering, volatility helps predict mechanical failures (MH1; MH2), while in hydrology, it informs understanding of streamflow variations (geophysics4; geophysics1; geophysics2; geophysics3).
@@
Figure 1: Figures 1(a), 1(b) and 1(c) display the S&P 500 index from 2001-01-01 to 2021-12-31, the daily total electricity demand in New York State from 2015-07-01 to 2024-06-30, and the daily bike rental counts via Capital Bikeshare by non-members in the Washington metropolitan area from 2012-01-01 to 2017-12-31, respectively. Figures 1(d), 1(e) and 1(f) show the corresponding mean-centered series: weekly log-returns for the S&P 500 index with the average log return subtracted, and series for electricity demand and bike rentals mean-centered using a smoothing spline. The S&P 500 index data are obtained using the quantmod package in R (quantmod) from Yahoo Finance (yahoo_finance_data), the electricity demand data from the New York Independent System Operator (NYISO) (nyis_electricity), and the bike rental data from Capital Bikeshare website (capital_bikeshare_data).
@@
Classical volatility models include AutoRegressive Conditional Heteroskedasticity (ARCH) (ARCH), Generalized ARCH (GARCH) (GARCH2), and Stochastic Volatility (SV) (SVop; SVop2; SVop3). They all make the stationarity assumption that the unconditional variance is constant. However, a growing body of econometric literature points to evolving volatility patterns in stock markets around the world (NS2; NS3; UK; taiwan; chinese; SVRW_w). Moreover, the stationarity assumption is often unrealistic in fields like climate science and epidemiology, due to climate change and seasonal or outbreak-driven dynamics (climate1; epidem3).
@@
Figure 1 illustrates evolving volatility patterns in three real-world datasets: weekly log-returns of the S&P 500 index, daily electricity demand in New York State, and daily bike rental counts via Capital Bikeshare. All three exhibit nonstationary volatility patterns. In Figures 1(a) and 1(d), persistent periods of high and low volatility reflect changing market conditions. In Figures 1(b), 1(e), 1(c) and 1(f), both the electricity and bike data show seasonal cycles in both mean and volatility, with elevated volatility during high-demand periods.
@@
Existing approaches address such patterns by incorporating hidden Markov switching (HMS) structure into existing SV and GARCH models (MSe1; MSe2; MS4; MSARCH; MSGarch; MSGARCH2; MS1; MSSV). However, HMS models suffer from three major limitations: they (1) require the specification of the number of regimes, which is typically unknown; (2) are computationally intensive, as pointed out in MSefficiency1 and MSefficiency2; and (3) impose a discrete regime-switching structure, which may not capture ever-evolving volatility dynamics. For example, Figure 1(f) reveals a seasonal increase in volatility for bike rentals, with the magnitude of these fluctuations growing over time. Specifically, volatility in 2012 is substantially lower than in 2017. Such a pattern is difficult to reconcile with a model that is based on a fixed number of regimes, as it reflects a continuous evolution of volatility rather than discrete shifts.
@@
We introduce the Adaptive Stochastic Volatility (ASV) model as an extension of the Random Walk Stochastic Volatility (RWSV) model (SVRW1; SVRW2). RWSV assumes that the log-variance of the observed process follows a Gaussian random walk with a constant variance. In contrast, ASV introduces two significant modifications to RWSV: it (1) incorporates a time-varying variance for the log-variance increments; and (2) adopts a global-local shrinkage prior for adaptability. For the shrinkage prior, we specifically highlight the Dynamic Shrinkage Processes (DSP) by dsp, as it provides a flexible framework for adaptive shrinkage in time series. To further improve robustness, we consider a nugget variant of ASV, which adds an additional Gaussian noise term to capture local fluctuations not explained by the smooth latent process.
@@
The standout feature of our proposed model is its local adaptability, effectively estimating volatility in the presence of both gradual and abrupt changes in the volatility process, with notable smoothness in between. It exhibits robustness against model misspecification, enabling accurate recovery of the latent volatility structure across diverse data-generating processes. In addition, ASV’s smooth volatility estimates provide clearer insights into the underlying patterns and trends. The model also naturally highlights periods of heightened volatility variation, offering additional interpretability in identifying structural shifts.
@@
The remainder of the paper is organized as follows. Section 2 introduces DSP, presents its theoretical properties, and describes the proposed ASV model. Section 3 outlines the data augmentation strategy for efficient Gibbs sampling, with full conditionals provided in Appendix B. Section 4 reports simulation results, and Section 5 presents empirical analyses on data sets shown in Figure 1. In section 6, the Bayesian Trend Filter with ASV (BTF-ASV), which jointly estimates the time-varying means and the variances, is proposed as an extension and applied to the sunspot data (sunspot).
@@ 2 Methodology
The goal of this section is to present the proposed ASV model, which captures evolving volatility with a global-local shrinkage prior. We begin with the RWSV model, the baseline formulation with a constant innovation variance. We then introduce an intermediate extension with a local-only shrinkage prior. We subsequently describe the DSP, a global-local shrinkage prior. Finally, we present the full ASV model and its nugget extension.
@@ 2.1 The Random Walk Stochastic Volatility Model
We begin by reviewing the RWSV model (SVRW1; SVRW2). Consider a zero mean process with $T$ observations, $\{y_{t}\}_{t=1}^{T}$, and its log-variance term $\{h_{t}\}_{t=1}^{T}$. In the SV framework, the observations $y_{t}$ follow a normal distribution whose log-variance $h_{t}$ evolves according to a first-order autoregression. RWSV is a special case of SV in which the autoregressive coefficient is fixed at 1. The model is defined as:
@@
In RWSV, $h_{t}$ is governed by a time-invariant variance term $\sigma_{h}^{2}$, which determines the degree of variation between successive log-variances. A large $\sigma_{h}^{2}$ implies high probability of significant changes in the $h_{t}$ process, whereas a small $\sigma_{h}^{2}$ indicates a high probability of minimal to no changes in $h_{t}$. In a frequentist framework, $\sigma_{h}$ is treated as a fixed parameter, often estimated via quasi-maximum likelihood (SVRW1; SVRW2). A more recent study by SVRW_w employs a Bayesian framework for estimation with a non-informative inverse-gamma prior for $\pi(\sigma^{2}_{h})$.
@@ 2.2 Stochastic Volatility with Local Only Shrinkage
RWSV may fit poorly when the time series exhibits both significant and minimal changes, as it assumes a time-invariant innovation variance $\sigma^{2}_{h}$. Ideally, we would prefer $\sigma^{2}_{h}$ to adapt over time, remaining small during stable periods and increasing during periods of abrupt changes in $h_{t}$. A natural approach to address this limitation is to introduce sparsity in the variance process. Under a frequentist framework, sparsity in $\Delta h_{t}$ may be attained by imposing a penalty, such as the $\ell_{1}$ penalty used in LASSO (LASSO). In a Bayesian setting, near-sparsity can be imposed via heavy-tailed priors on the variance parameter. Motivated by the Bayesian LASSO (BayesianLasso) and the scale-mixture Gaussian representation of the Laplace distribution (laplace_scale), we propose an intermediate model, RWSV with Bayesian LASSO (RWSV-BL):
$[\Delta h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma^{2}_{h,t})$ $[\sigma^{2}_{h,t}|\Lambda]\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Exp}\{(2\Lambda^{2})^{-1}\},$
where $\Lambda>0$ is a hyperparameter controlling the prior concentration. Following BayesianLasso, we place a Gamma prior on $\Lambda^{2}$ for full Bayesian inference. Under RWSV-BL, $\sigma^{2}_{h,t}$ induces shrinkage by encouraging most increments to be small while still allowing occasional large deviations.
@@ 2.3 Dynamic Shrinkage Processes and their Properties
Near-sparsity may also be achieved through global-local shrinkage priors. These priors introduce a global parameter that controls the overall degree of shrinkage across time and local parameters that allow individual components to deviate when supported by the data. Given a sequence of parameters $\{\Theta_{t}\}_{t=1}^{T}$, they take the following hierarchical form:
$\Theta_{t}\sim N(0,\tau^{2}\lambda_{t}^{2}),\qquad\tau\sim\pi(\tau),\quad\lambda_{t}\sim\pi(\lambda_{t}),$
where $\tau$ governs global shrinkage and $\lambda_{t}$ captures local flexibility. This class of priors encourages most parameters to be strongly regularized while permitting occasional large deviations, making it well suited for volatility estimation, where periods of stability are occasionally interrupted by abrupt shifts. This structure naturally extends the SV formulation by placing the prior on the $k$th-order difference of the log-variance process, $\{\Delta^{k}h_{t}\}_{t=1}^{T}$, so that its variance decomposes as $\sigma^{2}_{h,t}=\tau^{2}\lambda_{t}^{2}$.
Global-local shrinkage priors have been extensively studied in the context of high-dimensional Gaussian models, with notable examples including the horseshoe prior (horseshoe_theory), the horseshoe+ prior (horseshoe_P), and the triple-gamma prior (tgamma). For time-series modeling, several related dynamic shrinkage formulations have been proposed. dsptype1 and dsptype2 extend horseshoe-type priors using heavy-tailed $t$ distributions and fusion-based structures. These priors share the same underlying motivation as the DSP, combining global regularization with locally adaptive shrinkage for recovering the underlying temporal structure.
@@
Among these priors, DSP offers a particularly flexible formulation that generalizes several existing priors and naturally incorporates temporal dependence. DSP has shown strong versatility across various applications. abco integrate DSP into a Bayesian dynamic linear model to estimate change points and detect outliers. dspCount apply it to develop the negative binomial Bayesian trend filter (NB-BTF), a method for smoothing integer-valued time series. Additionally, dspDrift combine the Bayesian trend filter with DSP and a machine-learning-based regularization method to effectively distinguish micro-level drifts from macro-level shifts.
@@
A defining feature of DSP is its autoregressive structure on its log-variance process. It enables locally adaptive shrinkage through heavy-tailed innovations around a persistent mean process. Let $v_{t}:=\log(\sigma^{2}_{h,t})=\log(\tau^{2}\lambda_{t}^{2})$. $v_{t}\mid\mu,\phi\sim DSP(a,b,\mu,\phi)$ is defined by:
$v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t}$ $[\eta_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}Z(a,b,0,1),$ (3)
where $Z$-distribution (zdist) has density:
$f(z|a,b,\mu_{z},\sigma_{z})=(\sigma_{z}\mathrm{B}(a,b))^{-1}\bigg(\exp\bigg\{\frac{z-\mu_{z}}{\sigma_{z}}\bigg\}\bigg)^{a}\bigg(1+\exp\bigg\{\frac{z-\mu_{z}}{\sigma_{z}}\bigg\}\bigg)^{-(a+b)}.$
For fully Bayesian inference, we place priors on $\mu$ and $\phi$, and describe them in detail in Section 3.
@@
Under DSP, the global parameter is defined as $\tau=\exp\{\mu(1-\phi)/2\}$, and the local shrinkage parameter as $\lambda_{t}=\exp\{(\phi v_{t-1}+\eta_{t})/2\}$. The coefficient $|\phi|<1$ controls temporal dependence in local shrinkage. When $\phi$ is close to 1, the process exhibits strong positive dependence: strong shrinkage tends to be followed by continued strong shrinkage, and weak shrinkage by weak shrinkage. When $\phi$ is close to -1, the process favors alternation between strong and weak shrinkage over time.
@@
The DSP includes several well-known priors as special cases. When $a=b=1/2$ and the autoregressive coefficient $\phi=0$, it reduces to the horseshoe (HS) prior (horseshoe_theory), since $\lambda_{t}=\exp\{\eta_{t}/2\}\sim C^{+}(0,1)$. Keeping $a=b=1/2$ but assigning $\phi$ its own prior (e.g., Beta) introduces temporal dependence in shrinkage, yielding the Dynamic Horseshoe (DHS) prior (dsp). Other notable cases include the Strawderman–Berger prior when $a=1/2$ and $b=1$ (strawderman), the normal–exponential–gamma prior when $a=1$ and $b>2$ (griffin2005), and the Jeffreys prior with $a=b\rightarrow 0$ (figueiredo).
@@
While DSP has an appealing formulation and a clear connection to well-known priors, its theoretical properties remain incomplete. In particular, DSP distinguishes itself from other global-local shrinkage priors by introducing temporal dependence in the local shrinkage parameter $\lambda_{t}$. Due to the time dependence, the original work by dsp focuses on the conditional behavior, leaving the properties of the marginal distribution unexplored. We address this gap by deriving new results on the stationary distribution of DSP.
@@
Define $z_{\ell,t}:=\phi^{\ell}\eta_{t-\ell}$. The infinite-order moving average representation of $v_{t}$ is given by $v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t}=\mu+\sum_{\ell=0}^{\infty}z_{\ell,t}$, where $\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$. As shown in zdist, $z_{\ell,t}$ follows a scaled hyperbolic secant distribution, with $\mathbb{E}(z_{\ell,t}|\phi)=0$ and $\mathrm{Var}(z_{\ell,t}|\phi)=(|\phi|^{\ell}\pi)^{2}$. Let $z_{t}:=\sum_{\ell=0}^{\infty}z_{\ell,t}$ denote the aggregated process.
@@
The series $z_{t}$ converges almost surely if and only if $|\phi|<1.$
@@
This result directly follows from Kolmogorov’s three-series theorem. The variance $\sum_{\ell=0}^{\infty}\mathrm{Var}(z_{\ell,t}\mid\phi)$ converges if and only if $|\phi|<1$. The exact derivation is provided in Appendix A.1. ∎
@@
If $|\phi|<1$, then
$\mathbb{E}(z_{t}|\phi)=\mathbb{E}\bigg(\sum_{\ell=0}^{\infty}z_{\ell,t}|\phi\bigg)=\sum_{\ell=0}^{\infty}\mathbb{E}(z_{\ell,t}|\phi)=0.$
$\mathrm{Var}(z_{t}|\phi)=\mathrm{Var}\bigg(\sum_{\ell=0}^{\infty}z_{\ell,t}|\phi\bigg)=\sum_{\ell=0}^{\infty}\mathrm{Var}(z_{\ell,t}|\phi)=\sum_{\ell=0}^{\infty}\pi^{2}\phi^{2\ell}=\frac{\pi^{2}}{1-\phi^{2}}.$
The characteristic function of $z_{\ell,t}$ is $\mathrm{sech}(\pi\phi^{\ell}t)$. Thus, the characteristic function, $g(t)$ for $\sum_{\ell=0}^{\infty}z_{\ell,t}$ would be the infinite product of the characteristic function of $z_{\ell,t}$.
$g(t)=\prod_{\ell=0}^{\infty}\mathrm{sech}(\pi\phi^{\ell}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
Similarly, for the moment generating function $M(t)$ is given by
$M(t)=\prod_{\ell=0}^{\infty}\mathrm{sec}(\pi\phi^{\ell}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
@@
If $|\phi|=0.5$, $\eta_{t-\ell}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$, and $z_{\ell,t}:=\phi^{\ell}\eta_{t-\ell}$, then $z_{t}:=\sum_{\ell=0}^{\infty}z_{\ell,t}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\mathrm{Logistic}(0,2)$.